# CAST: Cross-Attentive Spatio-Temporal feature fusion for Deepfake detection

### v1

@@ 4.3 Evaluation Metrics:
To quantitatively assess the performance of our proposed deepfake detection model, we utilize several standard binary classification metrics, including Accuracy and AUC (Area Under the Curve). These metrics are computed based on the predicted labels for each video, where each video-level prediction is obtained by averaging the model’s sigmoid outputs across the 16 extracted frames and applying a threshold of 0.5.
Let $TP$ (True Positives) represent the number of fake videos correctly classified as fake, $TN$ (True Negatives) the number of real videos correctly classified as real, $FP$ (False Positives) the number of real videos incorrectly classified as fake, and $FN$ (False Negatives) the number of fake videos incorrectly classified as real. The evaluation metrics are formulated as follows:
@@ 5 Experimental results and discussion:
To comprehensively evaluate the performance and generalization capability of our proposed method, we conduct experiments across four key evaluation protocols: single-manipulation evaluation, intra-dataset testing, cross-dataset generalization, and multi-source manipulation robustness. These evaluations are performed primarily on the FaceForensics++ [1] dataset under high-quality compression settings, along with DFD [31] and Celeb-DF (v2) [2] datasets. We adopt standard classification metrics, including AUC and Accuracy, to ensure a consistent and fair comparison with existing state-of-the-art (SOTA) approaches.
@@ 5.1 Single Manipulation Evaluation
Table 2: Single-manipulation training evaluation on FF++ (HQ) in terms of AUC (%). Each model is trained on one specific manipulation subset and tested across all four subsets (DF, F2F, FS, NT).
Training Set Method DF F2F FS NT Avg D2-Fusion [10] 99.98 77.88 62.25 75.73 78.96 DCL [27] 99.98 77.13 61.01 75.01 78.28 RECCE [32] 99.65 70.66 74.29 67.34 70.76 DF ADAL [33] 99.68 79.41 73.61 81.92 83.65 WATCHER [34] 99.68 75.50 79.51 83.66 84.58 CAST-B5(Ours) 99.95 82.43 59.88 86.41 82.16 CAST-B0(Ours) 100.00 86.29 60.15 92.02 84.65 D2-Fusion [10] 89.50 99.86 62.47 75.23 81.76 DCL [27] 91.91 99.21 59.58 66.67 79.34 RECCE [32] 75.99 98.06 64.53 72.32 70.95 F2F ADAL [33] 90.32 99.17 69.49 73.13 83.03 WATCHER [34] 85.73 99.82 69.38 74.16 82.27 CAST-B5(Ours) 88.64 98.61 65.69 70.19 80.78 CAST-B0(Ours) 94.33 99.68 58.42 82.74 83.79 D2-Fusion [10] 77.50 69.76 99.92 58.45 76.41 DCL [27] 74.80 69.75 99.90 52.60 74.26 RECCE [32] 82.39 64.44 98.82 56.70 67.84 FS ADAL [33] 72.36 70.20 99.91 62.12 76.14 WATCHER [34] 84.52 75.06 99.95 58.81 79.58 CAST-B5(Ours) 58.13 74.71 99.99 53.61 71.61 CAST-B0(Ours) 59.97 83.39 99.29 51.11 73.44 D2-Fusion [10] 94.44 71.08 80.75 99.43 86.42 DCL [27] 91.23 79.31 52.13 98.97 80.41 RECCE [32] 78.83 80.89 63.70 93.63 74.47 NT ADAL [33] 90.94 63.28 78.47 99.28 82.99 WATCHER [34] 89.67 72.85 69.96 98.58 82.76 CAST-B5(Ours) 95.6 77.73 62.17 97.97 83.36 CAST-B0(Ours) 95.27 74.42 51.33 97.82 79.71
@@ 5.2 Intra-dataset evaluation on FF++
We evaluate the performance of our proposed models under the intra-dataset setting using the high-quality (HQ) subset of the FF++ dataset. Table 3 compares our EfficientNet-based models (CAST-B0 and CAST-B5) against several SOTA deepfake detection methods in terms of accuracy (ACC) and area under the ROC curve (AUC). Among existing methods, RECCE [32], MAT [15] and D2-Fusion [10] demonstrate strong performance in terms of AUC (%). Our proposed CACT-B5 model achieves the highest AUC of 99.49% as compared to these methods. Although CAST-B5 achieved 97.57% accuracy which is slightly lower than D2-Fusion who achieved 97.77%, it maintains a strong accuracy against MAT (97.60%). Furthermore, CACT-B5 outperforms other notable methods such as F3-Net [35] with 97.52%, GFF [36] with 96.87%, and RECCE with 97.06%, indicating a clear margin of improvement in accuracy and highlighting the effectiveness of our approach. The CAST-B0 model, though slightly lower in performance, still achieves competitive results with 96.89% accuracy and 98.98% AUC. This result highlights its efficacy as a lightweight alternative while retaining robust detection performance. The relatively small performance gap between CAST-B0 and CAST-B5 also suggests that the cross-attention mechanism contributes significantly to generalization and representation power, even when used with a lower-capacity CNN backbone. These results confirm the effectiveness of our CNN-Transformer architecture in capturing both spatial and temporal features for intra-dataset detection. The performance of CAST-B5, demonstrates impressive performance on FF++ (HQ), showing that our cross-attention design effectively bridges spatial cues from individual frames with sequence-level temporal context.
@@
Table 4: Cross-dataset evaluation on Celeb-DF v2 and DFD by training on FF++ (HQ), compared with other methods in terms of AUC (%).
Method Year AUC (%) DFD Celeb-DF Face X-ray [3] 2020 85.60 74.20 GFF [36] 2021 91.90 79.40 MAT [15] 2021 - 67.40 LTW [37] 2021 88.56 77.14 RECCE [32] 2022 - 68.71 DCL [27] 2022 91.66 82.30 SFDG [38] 2023 88.00 75.83 ID-MAM [39] 2025 - 72.04 CAST-B5 (Ours) 2025 92.07 74.92 CAST-B0 (Ours) 2025 93.31 76.98
@@ 5.3 Cross-dataset evaluation
To further validate the generalization ability of our proposed models, we perform cross-dataset evaluation by training on the high-quality (HQ) subset of FaceForensics++ (FF++) and testing on two challenging unseen datasets: Celeb-DF (v2) and DeepFakeDetection (DFD). This evaluation simulates realistic deployment conditions where the model encounters manipulation methods and identities that were not part of its training distribution. Table 4 presents the results in terms of AUC (%) for both datasets, compared with leading state-of-the-art methods.
Among prior works, DCL [27] demonstrates the highest overall performance on the Celeb-DF benchmark, achieving an AUC of 82.30%, followed by GFF [36] and LTW [37], which attain AUC scores of 79.40% and 77.14%, respectively. On the DFD dataset, both GFF and DCL exhibit strong performance, reporting AUC scores of 91.90% and 91.66% respectively. In contrast, our proposed CAST-B0 model achieves a superior AUC of 93.31% on DFD, indicating enhanced generalization capabilities compared to prior methods. On the more challenging Celeb-DF benchmark, characterized by high-quality and artifact-free manipulations, CAST-B0 achieves an AUC of 76.98%. While this performance is slightly below that of DCL, GFF, and LTW, it remains highly competitive. Furthermore, CAST-B0 surpasses other recent methods, including SFDG [38] (75.83%) and Face X-ray [3] (74.20%). The more complex CAST-B5 variant also yields competitive results, achieving AUC scores of 92.07% on DFD and 74.92% on Celeb-DF. Notably, it outperforms several state-of-the-art baselines such as Face X-ray [3] (74.20%), MAT [15] (67.40%), and ID-MAM [39] (72.04%).
In summary, the cross-dataset results confirm that our CAST-B0 model outperforms all SOTA methods for DFD and demonstrates competitive performance for Celeb-DF. The more complex CAST-B5 model achieved slightly lower AUC scores, but still outperformed majority of the SOTA methods. These results present an improved generalization across different deepfake generation methods and real-world video distributions.
@@ 5.4 Multi-source manipulation evaluation
To assess the robustness and generalization capability of our model across different manipulation types, we perform a multi-source evaluation on the FF++(HQ) dataset. In this setup, the model is trained on three manipulation methods and tested on the remaining unseen method. This cross-manipulation setting allows us to evaluate how well the model generalizes to novel forgery techniques not encountered during training. Table 5 presents the AUC (%) performance for several state-of-the-art models on three different manipulation types: Deepfakes (DF), Face2Face (F2F), and NeuralTextures (NT). Our CAST-B0 model demonstrates superior generalization across all three evaluation settings. On the DF split, CAST-B0 achieves an AUC of 97.04%, outperforming recent models like Implicit [40] (95.03%) and D2-Fusion [10] (95.34%). For the F2F subset, CAST-B0 again leads with an AUC of 87.50%, showing notable improvement over MAT [15] (78.52%). The most significant gain is observed in the NT split, where CAST-B0 achieves a substantial AUC of 84.50%, outperforming D2-Fusion by over 4% and showing a 16.3% improvement over MAT.
These results highlight the effectiveness of our CAST architecture in learning transferable representations that generalize across different manipulation methods. The consistent gains across all manipulation types highlight the model’s strong domain adaptability and robustness.
@@ 5.5 Ablation Studies
To assess the contribution of each key module in our proposed CAST framework, we conducted a comprehensive ablation study. All experiments are performed using the FF++(HQ) dataset for training, with both intra-dataset evaluation and cross-dataset generalization testing using Celeb-DF and DFD.
@@ 5.5.1 Impact of Cross-Attention Mechanism
To evaluate the influence of the cross-attention mechanism, we constructed a baseline CNN-Transformer model by removing the cross-attention block from the CAST model. In this configuration, temporal features derived from the transformer encoder using frame-wise CNN features were directly used for final classification without explicit spatial-temporal interaction. From results in Table 6, we observe that the absence of cross-attention led to a noticeable drop in cross-dataset generalization, with AUC scores decreasing to 66.18 on Celeb-DF and 90.10 on DFD both substantially lower than the CAST model. Although intra-dataset performance on FF++ remained relatively high (98.12 %), the model’s limited robustness across datasets underscores the necessity of cross-attention in aligning spatial and temporal representations. This mechanism enables the model to capture nuanced manipulation traces, thereby enhancing the consistency and reliability of detection under diverse real-world conditions.
@@ 5.5.2 Evaluating Cross-Attention against Decoupled Attention Fusion
To assess the contribution of direct spatial-temporal interaction, we replace the proposed cross-attention module with independent self-attention layers applied to the spatial and temporal feature streams, followed by feature fusion via concatenation. As shown in Table 6, this decoupled configuration results in an intra-dataset AUC of 96.38% on FF++ (HQ), notably lower than the CAST model. More critically, its performance on cross-dataset evaluations dropped to 71.41% on Celeb-DF and 87.24% on DFD, indicating diminished generalization. These findings suggest that self-attended spatial and temporal features, when combined without explicit interaction, lack the capacity to model complex manipulation artifacts such as asynchronous facial dynamics or localized geometric inconsistencies. In contrast, cross-attention provides a structured mechanism for aligning spatial and temporal representations, enabling more robust detection across varied data distributions.
Fig. 4 qualitatively supports the ablation results by comparing attention maps across manipulation types. Without cross-attention, the model exhibits diffuse and unfocused activations, while the separate self-attention setup leads to fragmented responses. In contrast, our cross-attention mechanism where temporal features attend over spatial representations produces sharper and more localized activations around manipulated regions. This demonstrates its ability to integrate spatial context dynamically, resulting in more interpretable and generalizable representations.
@@ 5.5.3 Effect of Cross-Attention Directionality
To examine the influence of attention directionality within the cross-attention module, we modified the original configuration by reversing the query and key-value roles by assigning spatial tokens as queries and temporal tokens as key-value inputs. As shown in Table 6, this reversed setup yielded an AUC of 97.93% on FF++ (HQ), which is marginally lower than the CAST model. However, a more pronounced drop was observed in cross-dataset settings, with performance decreasing to 71.25% on Celeb-DF, although DFD reduced by a small margin to 90.41%. These outcomes suggest that leveraging temporal tokens as queries to guide spatial feature refinement is more effective, likely due to temporal dynamics offering a stronger prior for identifying manipulation-relevant spatial regions. This directional design not only improves cross-domain robustness but also aids in the interpretability of the model by aligning temporal cues with spatial inconsistencies.

### v2

@@ 4.3 Evaluation Metrics:
To quantitatively assess the performance of our proposed deepfake detection model, we utilize several standard binary classification metrics, including Accuracy and Area Under the Curve (AUC). These metrics are computed based on the predicted labels for each video, where each video-level prediction is obtained by averaging the model’s sigmoid outputs across the 16 extracted frames and applying a threshold of 0.5.
Let True Positives ($TP$) represent the number of fake videos correctly classified as fake, True Negatives ($TN$) the number of real videos correctly classified as real, False Positives ($FP$) the number of real videos incorrectly classified as fake, and False Negatives ($FN$) the number of fake videos incorrectly classified as real. The evaluation metrics are formulated as follows:
@@ 5 Experimental results and discussion:
To comprehensively evaluate the performance and generalization capability of our proposed method, we conduct experiments across four key evaluation protocols: single-manipulation evaluation, intra-dataset testing, cross-dataset generalization, and multi-source manipulation assessment. These evaluations are performed primarily on the FaceForensics++ [33] dataset under high-quality compression settings, along with DFD [12], DFDC [10] and Celeb-DF (v2) [24] datasets. We adopt standard classification metrics, including AUC and Accuracy, to ensure a consistent and fair comparison with existing state-of-the-art (SOTA) approaches.
@@ 5.1 Single Manipulation Evaluation
Table 2: Single-manipulation training evaluation on FF++ (HQ) in terms of AUC (%). Each model is trained on one specific manipulation subset and tested across all four subsets (DF, F2F, FS, NT). Intra-manipulation results are highlighted in gray. EfficientNet-B0 and EfficientNet-B5 as CNN backbone variants are considered for comparison.
Training Set Method Year DF F2F FS NT Avg D2-Fusion [31] 2025 99.98 77.88 62.25 75.73 78.96 DCL [38] 2022 99.98 77.13 61.01 75.01 78.28 RECCE [3] 2022 99.65 70.66 74.29 67.34 70.76 DF ADAL [23] 2023 99.68 79.41 73.61 81.92 83.65 WATCHER [44] 2024 99.68 75.50 79.51 83.66 84.58 CAST-B5(Ours) 2025 99.95 82.43 59.88 86.41 82.16 CAST-B0(Ours) 2025 100.00 86.29 60.15 92.02 84.65 D2-Fusion [31] 2025 89.50 99.86 62.47 75.23 81.76 DCL [38] 2022 91.91 99.21 59.58 66.67 79.34 RECCE [3] 2022 75.99 98.06 64.53 72.32 70.95 F2F ADAL [23] 2023 90.32 99.17 69.49 73.13 83.03 WATCHER [44] 2024 85.73 99.82 69.38 74.16 82.27 CAST-B5(Ours) 2025 88.64 98.61 65.69 70.19 80.78 CAST-B0(Ours) 2025 94.33 99.68 58.42 82.74 83.79 D2-Fusion [31] 2025 77.50 69.76 99.92 58.45 76.41 DCL [38] 2022 74.80 69.75 99.90 52.60 74.26 RECCE [3] 2022 82.39 64.44 98.82 56.70 67.84 FS ADAL [23] 2023 72.36 70.20 99.91 62.12 76.14 WATCHER [44] 2024 84.52 75.06 99.95 58.81 79.58 CAST-B5(Ours) 2025 58.13 74.71 99.99 53.61 71.61 CAST-B0(Ours) 2025 59.97 83.39 99.29 51.11 73.44 D2-Fusion [31] 2025 94.44 71.08 80.75 99.43 86.42 DCL [38] 2022 91.23 79.31 52.13 98.97 80.41 RECCE [3] 2022 78.83 80.89 63.70 93.63 74.47 NT ADAL [23] 2023 90.94 63.28 78.47 99.28 82.99 WATCHER [44] 2024 89.67 72.85 69.96 98.58 82.76 CAST-B5(Ours) 2025 95.6 77.73 62.17 97.97 83.36 CAST-B0(Ours) 2025 95.27 74.42 51.33 97.82 79.71
@@ 5.2 Intra-dataset evaluation on FF++
We evaluate the performance of our proposed models under the intra-dataset setting using the high-quality (HQ) subset of the FF++ dataset. Table 3 compares our EfficientNet-based models (CAST-B0 and CAST-B5) against several SOTA deepfake detection methods in terms of accuracy (ACC) and area under the ROC curve (AUC). Among existing methods, RECCE [3], MAT [51] and D2-Fusion [31] demonstrate strong performance in terms of AUC (%). Our proposed CAST-B5 model achieves the highest AUC of 99.49% as compared to these methods. Although CAST-B5 achieved 97.57% accuracy which is slightly lower than D2-Fusion who achieved 97.77%, it maintains a strong accuracy against MAT (97.60%). Furthermore, CAST-B5 outperforms other notable methods such as F3-Net [30] with 97.52%, GFF [26] with 96.87%, and RECCE with 97.06%, indicating a clear margin of improvement in accuracy and highlighting the effectiveness of our approach. The CAST-B0 model, though slightly lower in performance, still achieves competitive results with 96.89% accuracy and 98.98% AUC. This result highlights its efficacy as a lightweight alternative while retaining robust detection performance. The relatively small performance gap between CAST-B0 and CAST-B5 also suggests that the cross-attention mechanism contributes significantly to generalization and representation power, even when used with a lower-capacity CNN backbone. These results confirm the effectiveness of our CNN-Transformer architecture in capturing both spatial and temporal features for intra-dataset detection. The performance of CAST-B5, demonstrates impressive performance on FF++ (HQ), showing that our cross-attention design effectively bridges spatial cues from individual frames with sequence-level temporal context.
@@
Table 4: Cross-dataset evaluation on Celeb-DF v2, DFD and DFDC by training on FF++ (HQ), compared with other methods in terms of AUC (%).
Method Year AUC (%) DFD Celeb-DF DFDC Face X-ray [22] 2020 85.60 74.20 70.00 GFF [26] 2021 91.90 79.40 79.70 MAT [51] 2021 - 67.40 - LTW [37] 2021 88.56 77.14 69.00 RECCE [3] 2022 - 68.71 69.06 DCL [38] 2022 91.66 82.30 76.71 SFDG [45] 2023 88.00 75.83 73.64 MeST-Former [25] 2024 72.16 76.33 - ID-MAM [34] 2025 - 72.04 - CAST-B5 (Ours) 2025 92.07 74.92 78.39 CAST-B0 (Ours) 2025 93.31 76.98 81.25
@@ 5.3 Cross-dataset evaluation
To further validate the generalization ability of our proposed models, we perform cross-dataset evaluation by training on the high-quality (HQ) subset of FaceForensics++ (FF++) and testing on three challenging unseen datasets: Celeb-DF (v2), DeepFakeDetection (DFD) and Deepfake Detection Challenge (DFDC). This evaluation simulates realistic deployment conditions where the model encounters manipulation methods and identities that were not part of its training distribution. Table 4 presents the results in terms of AUC (%) for both datasets, compared with leading state-of-the-art methods.
Among prior works, DCL [38] demonstrates the highest overall performance on the Celeb-DF benchmark, achieving an AUC of 82.30%, followed by GFF [26] and LTW [37], which attain AUC scores of 79.40% and 77.14%, respectively. On the DFD dataset, both GFF and DCL exhibit strong performance, reporting AUC scores of 91.90% and 91.66% respectively. In contrast, our proposed CAST-B0 model achieves a superior AUC of 93.31% on DFD, indicating enhanced generalization capabilities compared to prior methods. On the more challenging Celeb-DF benchmark, characterized by high-quality and artifact-free manipulations, CAST-B0 achieves an AUC of 76.98%. While this performance is slightly below that of DCL, GFF, and LTW, it remains highly competitive. Furthermore, CAST-B0 surpasses other recent methods, including SFDG [45] (75.83%), Face X-ray [22] (74.20%) and MeST-Former [25] (76.33%). On DFDC, GFF and DCL achieve AUC scores of 79.70% and 76.71% respectively, highlighting the increased difficulty of this dataset due to its diverse manipulations and realistic scenarios. Comparatively, our proposed CAST-B0 model achieves a superior AUC of 81.25% on DFDC, outperforming all prior methods on this benchmark. The more complex CAST-B5 variant also yields competitive results, achieving AUC scores of 92.07% on DFD, 74.92% on Celeb-DF and 78.39% on DFDC. Notably, it outperforms several state-of-the-art baselines such as Face X-ray [22] (74.20%), MAT [51] (67.40%), and ID-MAM [34] (72.04%).
In summary, the cross-dataset results confirm that our CAST-B0 model outperforms all SOTA methods on DFDC and DFD, whle demonstrating competitive performance on Celeb-DF. The CAST-B5 model achieved slightly lower AUC scores, but still outperformed the majority SOTA methods. These results present an improved generalization across different deepfake generation methods and datasets.
@@ 5.4 Multi-source manipulation evaluation
To assess the generalization capability of our model across different manipulation types, we perform a multi-source evaluation on the FF++(HQ) dataset. In this setup, the model is trained on three manipulation methods and tested on the remaining unseen method. This cross-manipulation setting allows us to evaluate how well the model generalizes to novel forgery techniques not encountered during training. Table 5 presents the AUC (%) performance for several state-of-the-art models on three different manipulation types: Deepfakes (DF), Face2Face (F2F), and NeuralTextures (NT). Our CAST-B0 model demonstrates superior generalization across all three evaluation settings. On the DF split, CAST-B0 achieves an AUC of 97.04%, outperforming recent models like Implicit [16] (95.03%) and D2-Fusion [31] (95.34%). For the F2F subset, CAST-B0 again leads with an AUC of 87.50%, showing notable improvement over MAT [51] (78.52%). The most significant gain is observed in the NT split, where CAST-B0 achieves a substantial AUC of 84.50%, outperforming D2-Fusion by over 4% and showing a 16.3% improvement over MAT.
These results highlight the effectiveness of our CAST architecture in learning transferable representations that generalize across different manipulation methods. The consistent gains across all manipulation types highlight the model’s strong domain adaptability.
@@
Model Variant AUC (%) FF++ (HQ) Celeb-DF DFD DFDC No Cross-Attn (CNN + Transformer) 98.12 66.18 90.10 69.26 Self-Attn (Spatial + Temporal sep.) 96.38 71.41 87.24 66.75 Cross-Attn reversed (Spatial $\rightarrow$ Temporal) 97.93 71.25 90.41 71.50 Multi-Scale Integration (CNN Layers) 98.04 72.33 88.64 72.06 No projection layer 98.60 68.14 88.49 74.19 CAST-B0 98.98 76.98 93.31 81.25
@@ 5.5 Ablation Studies
To assess the contribution of each key module in our proposed CAST framework, we conducted a comprehensive ablation study. All experiments are performed using the FF++(HQ) dataset for training, with both intra-dataset evaluation and cross-dataset generalization testing using Celeb-DF, DFD and DFDC.
@@ 5.5.1 Impact of Cross-Attention Mechanism
To evaluate the influence of the cross-attention mechanism, we constructed a baseline CNN-Transformer model by removing the cross-attention block from the CAST model. In this configuration, temporal features derived from the transformer encoder using frame-wise CNN features were directly used for final classification without explicit spatial-temporal interaction. From results in Table 6, we observe that the absence of cross-attention led to a noticeable drop in cross-dataset generalization, with AUC scores decreasing to 66.18% on Celeb-DF, 90.10% on DFD and 69.26% on DFDC, all substantially lower than the CAST model. Although intra-dataset performance on FF++ remained relatively high (98.12 %), the model’s limited strength across datasets underscores the necessity of cross-attention in aligning spatial and temporal representations. This mechanism enables the model to capture nuanced manipulation traces, thereby enhancing the consistency and reliability of detection under diverse real-world conditions.
@@ 5.5.2 Evaluating Cross-Attention against Decoupled Attention Fusion
To assess the contribution of direct spatial-temporal interaction, we replace the proposed cross-attention module with independent self-attention layers applied to the spatial and temporal feature streams, followed by feature fusion via concatenation. As shown in Table 6, this decoupled configuration results in an intra-dataset AUC of 96.38% on FF++ (HQ), notably lower than the CAST model. More critically, its performance on cross-dataset evaluations dropped to 71.41% on Celeb-DF, 87.24% on DFD and 66.75% on DFDC, indicating diminished generalization. These findings suggest that self-attended spatial and temporal features, when combined without explicit interaction, lack the capacity to model complex manipulation artifacts such as asynchronous facial dynamics or localized geometric inconsistencies. In contrast, cross-attention provides a structured mechanism for aligning spatial and temporal representations, enabling more robust detection across varied data distributions.
Fig. 4 qualitatively supports the ablation results by comparing attention maps across manipulation types. Without cross-attention, the model exhibits diffuse and unfocused activations, while the separate self-attention setup leads to fragmented responses. In contrast, our cross-attention mechanism, where temporal features attend over spatial representations, produces sharper and more localized activations around manipulated regions. This demonstrates its ability to integrate spatial context dynamically, resulting in more interpretable and generalizable representations.
@@ 5.5.3 Effect of Cross-Attention Directionality
To examine the influence of attention directionality within the cross-attention module, we modified the original configuration by reversing the query and key-value roles by assigning spatial tokens as queries and temporal tokens as key-value inputs. As shown in Table 6, this reversed setup yielded an AUC of 97.93% on FF++ (HQ), which is marginally lower than the CAST model. However, a more pronounced drop was observed in cross-dataset settings, with performance decreasing to 71.25% on Celeb-DF and 71.50% on DFDC, although DFD reduced by a small margin to 90.41%. These outcomes suggest that leveraging temporal tokens as queries to guide spatial feature refinement is more effective, likely due to temporal dynamics offering a stronger prior for identifying manipulation-relevant spatial regions. This directional design not only improves cross-domain performance but also aids in the interpretability of the model by aligning temporal cues with spatial inconsistencies.
5.5.4 Effect of Multi-Scale Feature Integration
To analyze whether incorporating hierarchical spatial representations benefits the CAST architecture, we constructed a multi-scale variant using the same EfficientNet-B0 backbone. Intermediate feature maps from multiple stages of the backbone were extracted, concatenated and projected to the transformer embedding space for temporal modeling. These fused multi-scale embeddings were then passed as the key and value parameters of the cross-attention module, thereby isolating the effect of spatial scale aggregation without altering the backbone family or capacity class.
However, the inclusion of multi-scale features from CNN layers resulted in a consistent decline in performance compared to the baseline CAST-B0 model. The multi-scale variant achieved AUC scores of 98.04% on FF++ (HQ), 72.33% on Celeb-DF, 88.64% on DFD and 72.06% on DFDC, all notably lower than the corresponding CAST-B0 results. This degradation appears to stem from the representational mismatch introduced by combining heterogeneous low-level and high-level features into a single spatial embedding, which weakens the alignment between spatial frame descriptors and temporally encoded tokens. Since CAST relies on clean, semantically coherent spatial features for effective cross-attention, introducing such heterogeneous mixtures disrupts spatial–temporal correspondence and ultimately hampers the fusion process.
These results indicate that, within the CAST framework, relying on a single high-level semantic representation extracted from the CNN layer provides more stable and discriminative spatial cues than aggregating multi-scale features, reaffirming the importance of consistent spatial embedding quality.